# I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses

### v2

@@ 5 Investigation of the proposed hypotheses
Table 2: Performance comparison of models trained on data constructed by different methods across various tasks. Data points are labeled as low performance when their accuracy is more than 15% lower relative to the highest accuracy achieved in the same dataset using the same model.
@@ 5.1 Understand the impact of “familiarity”
Our observation 1 suggests that using GPT-4 generated data leads to better training effectiveness than using human-annotated ground truth. Our second hypothesis posits that this enhanced learning by LLMs on LLMs responses could be attributed to familiarity with the data. It is a common belief that GPT-4 often includes more detailed content than human annotators, which could contribute to improved training outcomes. However, we question whether the superior results from GPT-4 are solely due to the addition of more details, or if familiarity plays a significant role as well.
@@
To explore whether familiarity is a crucial factor, we designed the following experiment. We employed different methods to have GPT-4 generate data in a detailed, step-by-step manner. For each dataset, we established a control group: one group involved GPT-4 transforming ground truth into its own detailed style, while the other had GPT-4 mimic the style of the human-labeled ground truth, adding detail to build upon it in a step-by-step style. As shown in Table 2, in the vast majority of cases, the training effectiveness of the first group, which used GPT-4’s own style, was superior to that of the second group, which followed the human-annotated style. In addition, it is noteworthy that using the Mistral 7B model for training on math algebra, converting ground truth into step-by-step answers was even detrimental. Both of the step by step groundtruth transformation methods achieves accuracy below 24% on Math Algebra, much lower than 30.2% achieved by ’GPT-4 answer directly’(Gpt-4 answer directly represents employing GPT-4 to generate responses using only the questions from the training datasets. For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation.) We speculate that one possible reason for this is that adding ’more than enough details’ may sometimes complicate the model’s reasoning process. We have included additional experiments involving various GPT-4 generated variants in the appendixA, which generally conform to the patterns described above. For clarity and conciseness, Table 2 only includes 2 examples of the step-by-step variants.
@@
After preliminary validation of our hypotheses, we made another intriguing observation: According to Table 2, the ’GPT-4 answer directly’ dataset, characterized by a shorter average token length of target labels compared to both ’step by step’ and ’step by step transformation of GT’ datasets, suggests fewer details in its content. Despite this, the performance of the data directly generated by GPT-4 frequently ranks among the best of all GPT-4 answer variants. Meanwhile, other data generation methods exhibited significantly lower performance than GPT-4’s direct answers on at least two tasks, as highlighted in red in Table 2. This suggests that, given a problem, the data directly generated by GPT-4 might yield the good training results. Artificially defined prompts that add information to the ground truth may lead to the generation of target labels that do not conform to the LLM’s linguistic or logical preferences.
Experiment 2: Direct measuring the impact of familiarity
@@
GPT-4 answer with lower perplexity Mistral 0.597 2.35 0.301 2.20 0.722 6.35
@@
GPT-4 answer with lower perplexity LLaMA 0.428 2.30 0.150 2.59 0.656 5.33
@@
Table 3: Comparison between the GPT-4 answer with higher perplexity Vs. lower perplexity
This section introduces a method for assessing the influence of style ’familiarity’. We direct GPT-4 to produce several varient of responses to the same query, grouping them into two distinct datasets based on their perplexity levels—high and low. We then evaluate and compare the learning outcomes using these two sets. One concern of this experiment is that GPT-4 may produce answers with different level of details. To eliminate this factor, we first let GPT-4 to generate one answer and then explicitly direct it to paraphrase the generated answer. In this way, we can create responses that vary in perplexity but maintain the same semantic meaning.
After training on those two sets of training data, we examine their performance, shown in Table 3. As seen, training on the higher-perplexity training set consistently shows slightly lower performance than the lower perplexity counterpart, particularly noticeable in the in-domain performance on two math datasets. For example, in the algebra dataset from Mistral, the lower perplexity set achieves a 30% accuracy rate, while the higher perplexity set only reached 23%. Since they are only differed in the style of the text, this result demonstrated that the familiarity of LLMs with the stylistic aspects of target labels significantly affects training outcomes.
@@
Discussion: Why are LLMs often familiar with responses generated from other LLMs? The familiarity observed when training on data generated by GPT-4 might stems from the overlap in training data used during the pre-training stages of these models. Many LLMs, including GPT-4, share similar pre-training data, developing an understanding of specific output styles and content.
@@ 5.2 Achieving high familiarity by exploring the target LLM generated responses
Math Commensense Reasoning NLI
Method Model Type GSM8K Math Algebra ECQA ESNLI
Groundtruth Mistral 0.434 0.206 0.68 0.814
Correct predicted sample + groundtruth 0.491 0.227 0.578 0.557
correct preidcted sample only 0.449 0.257 0.586 0.698
Style Transfered Groundtruth 0.51 19.6 0.696 0.845
Table 4: Performance Comparison when not using GPT-4. For the style-transferred ground truth dataset, only 73.7% and 24% of the training data are used for training, as explained in Section5.2
In the experiments described above, we validated two hypotheses. However, all the studies mentioned use GPT-4 to generate responses, chosen for its advanced capabilities to directly produce correct answers and reliably follow diverse instructions. However, there is a concern regarding the "familiarity" factor: would this argument still hold without the involvement of an advanced LLM? If valid, the argument should apply to some extent even without using advanced LLMs.
This section seeks to develop a method that enhances the “familiarity” of responses without requiring more advanced large language models (LLMs) than the target LLM.
We included ESNLICamburu et al. (2018) as an additional dataset to better compare different data construction methods. The e-SNLI dataset enhances the SNLIBowman et al. (2015) dataset by adding human-annotated explanations to each instance. These annotations offer insights into the reasoning behind the labels of entailment, contradiction, or neutral, improving model interpretability and aiding in the development of more robust inference systems.
Use the same model for train and groundtruth style transfer We devised a method to rewrite the ground truth data in a style that resembles language model-generated responses using the Mistral-7B-Instruct-V2 model. The experiment results are shown in Table 4. Initially, we posed questions from the training dataset and recorded the answers that were correctly predicted by Mistral. We documented two such correct predictions. These two correct responses were then used as in-context examples to guide Mistral on how to rewrite ground truth data. Specifically, we used the two recorded predictions as example target labels and their corresponding questions as example questions, while the actual ground truth served as the example ground truth. The in-context prompt informed Mistral that its task was to rewrite the ground truth in its own styles.
@@
We noticed that even when provided with the ground truth, the Mistral model struggles to successfully rewrite answers for mathematical tasks. Therefore, we ran the rewritten math answers through our evaluation script and found that only 73.7% of the GSM8K answers and 24% of the math algebra answers passed the script. For math tasks, we only used the answers that passed the script for training. In the ECQA and ESNLI tasks, we allowed the Mistral model to rewrite the ground truth once and directly used all the transformed ground truth answers for training.
@@
This experiment was conducted solely with Mistral and not with the Llama model, due to our observations that Llama2-13b-chat lacks sufficient in-context following capabilities required for successful ground truth style transformation. We show a failure example in AppendixD. As a comparative experiment, we created the ’correct predicted samples only’ dataset. For each sample in this dataset, we had the Mistral model perform zero-shot generation 5 times, and using an evaluation filter, we randomly selected one correct prediction that passed the filter to be paired with the original input as the target label. If the model failed to generate a correct answer in all 5 attempts, we did not include that data in the training dataset. The construction method for the ’correct predicted samples’ and ’correct predicted samples only’ datasets is similar. The only difference is that if the sample fails to produce a correct answer, we include the ground truth in the dataset as the correct answer.
@@
As shown in Table 4, the Mistral model trained with style-transferred ground truth data outperformed the model trained with original ground truth data and other methods in all cases except for the Math Algebra dataset, where only 24% of the training data was used. To fairly compare the training effectiveness of the ’style transferred ground truth’ and the original ground truth, we recorded the IDs of the data used by the ’style transferred ground truth’. We identified a corresponding set of 240 original ground truth data points for comparison. We found that training with only these 240 data points resulted in a 17.4% accuracy using the original ground truth, which is lower than the 19.6% achieved by the ’style transferred ground truth’. This demonstrates that data processed using the ’ground truth style transfer’ method yields better training outcomes than data labeled by humans.
@@
Interestingly, on the ECQA and ESNLI datasets, models trained with correct answers generated by Mistral in a zero-shot fashion performed worse than those trained directly with ground truth. By reading the ’correct predicted samples only’ dataset, we found that many instances classified as correct by the evaluation filter did not result from sound reasoning by the model. Although the final answers were technically correct, the underlying rationales were incorrect. Therefore, the effectiveness of the self-training method depends significantly on whether the evaluation filter used to select correct predictions can truly identify accurate responses.
@@ 6 A hybrid approach: creating responses using “minimal change” principal
We call this approach “minimal change” and its construction process is as follows: For a given question, we first request a response from the target LLM, such as Llama2 or Mistral. We then instruct GPT-4 to copy the correct portions of the prediction and revise any erroneous parts of the response. When correcting errors, GPT-4 is required to make minimal changes to accurately adjust the answer  [^1] Please note that achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal. Note that this construction process ensures that the target LLM is “familiar” with the response as the response is largely generated from itself. The “minimal change” principal, on the other hand, minimizes the influence from the GPT-4.
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt and Method’ section of the Appendix C, which includes two to three examples of minimal changes along with explanations for each modification.
Method Model Type GSM8K Math Algebra ECQA MBPP HumanEval Avg Perplexity
Groundtruth Mistral 0.434 0.206 0.68 0.276 0.323 15.03
GPT-4 Answer Directly 0.597 0.302 0.722 0.354 0.365 3.81
Minimum Change on Mistral Predictions 0.562 0.314 0.699 0.354 0.409 2.47
Minimum Change on LLaMA predictions 0.547 0.297 0.709 0.364 0.427 3.51
Groundtruth Llama 0.364 0.126 0.631 0.2 0.146 6.18
GPT-4 Answer Directly 0.428 0.150 0.656 0.2 0.158 3.58
Minimum Change on LLaMA predictions 0.433 0.166 0.649 0.2 0.213 2.75
Minimum Change on Mistral Predictions 0.402 0.161 0.647 0.218 0.183 3.32
Table 5: Comparing the experimental results of GPT4, groundtruth and minimum change
Results from our ’Minimum Change’ experiments are presented in Table 3. The data shows consistently lower perplexity compared to ground truth data across all datasets, indicating a high level of model familiarity with LLM-generated data. Furthermore, the performance across various setups was not only similar (with in-domain performance differences within 5% on most datasets) but also comparable to models trained with direct GPT-4 responses. Although the data was minimally modified based solely on the outputs from Mistral/LLaMA, theoretically, the logic and quality of the modified answers, as well as the level of detail, should be inferior to answers directly generated by GPT-4. Surprisingly, on the HumanEval and math algebra tasks, the performance of models trained with minimally changed data even surpassed those trained with data generated by GPT-4. On other datasets, the performance of the minimum change data was comparable to that of GPT-4. This underscores the effectiveness of training LLMs on data distributions familiar to them, demonstrating that minimal changes to maintain the original structure can be as effective as more detailed alterations. For more information about this experiment, please refer to the Appendix. B
@@ 7 Related Works
There are many studies related to using LLMs to generate training data for training models.
Self-Training: Several works utilize a model’s own predictions for self-training. For instance, STAR Zelikman et al. (2022) and REST Gulcehre et al. (2023) generates a dataset through sample production from the LLMs, subsequently utilizing these samples to enhance the LLMs via training. RESTem Singh et al. (2023) enhances model performance by using initial predictions, filtering for accuracy, and retraining the model with correct predictions. This iterative process improves the model’s accuracy over multiple cycles.
Data Augmentation: Numerous studies have investigated the use of synthetic data primarily for data augmentation purposes. In scenarios where sample sizes are insufficient, these studies generate additional training data to augment existing datasets (usually when the original training data is not adequate) Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Agrawal et al. (2023); Piedboeuf and Langlais (2023), or for generating extensive training data when high-quality human-annotated instruction tuning data is difficult to obtain Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). In cases where training data quality is low, many studies improve model training by providing more rationales Zhang et al. (2024); Kang et al. (2023); Li et al. (2022) or step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023) to small models.
Our work distinguishes itself from existing studies by focusing on the role of familiarity in training. We explore the effectiveness of synthetic data from the perspective of familiarity, providing a theoretical foundation that substantiates the efficacy of past methods using synthetic data. Our finding offers a nuanced perspective that underscores the importance of dataset familiarity in enhancing model performance.

### v3

@@ 4 Investigation of the hypotheses
Table 2: Performance comparison of models trained on data constructed using different methods. $n_{\text{train}}=1000$. Data points are labeled as low performance when accuracy is more than 15% below the highest accuracy on the same dataset using the same model. See Table 6 and Table 7 for additional experiments with GPT-4 and Claude 3.5.
@@ 4.1 Understand the impact of “familiarity”
Our observation 1 suggests that using advanced LLMs generated data leads to better training effectiveness than using human-annotated ground truth. Our second hypothesis posits that this enhanced learning by LLMs on LLMs responses could be attributed to familiarity with the data. It is a common belief that LLMs often includes more detailed content than human annotators, which could contribute to improved training outcomes. However, we question whether the superior results from advanced LLMs are solely due to the addition of more details, or if familiarity plays a significant role as well.
@@
To investigate whether the superior performance is due to the additional detail provided by advanced LLMs compared to humans, we designed the following experiment. We employed different methods to have GPT-4 generate data in a detailed, step-by-step manner. For each dataset, we established a control group: one group involved GPT-4 transforming ground truth into its own detailed style, while the other had GPT-4 mimic the style of the human-labeled ground truth, adding detail to build upon it in a step-by-step style. As shown in Table 2, in the vast majority of cases, the training effectiveness of the first group, which used GPT-4’s own style, was superior to that of the second group, which followed the human-annotated style. In addition, it is noteworthy that using the Mistral 7B model for training on math algebra, converting ground truth into step-by-step answers was even detrimental. Both of the step by step groundtruth transformation methods achieves accuracy below 24% on Math Algebra, much lower than 30.2% achieved by ‘GPT-4 answer directly’(Gpt-4 answer directly represents employing GPT-4 to generate responses using only the questions from the training datasets. For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation.) We speculate that one possible reason for this is that adding ‘more than enough details’ may sometimes complicate the model’s reasoning process. We have included additional experiments in the appendix A, involving various GPT-4 (refer to Table 6) and Claude 3.5 (refer to Table 7) generated variants with different token lengths (shorter or longer), which generally conform to the patterns described above. For clarity and conciseness, Table 2 only includes 2 examples of the step-by-step variants.
@@
After preliminary validation of our hypotheses, we made another intriguing observation: According to Table 2, the ‘GPT-4 answer directly’ dataset, characterized by a shorter average token length of target responses compared to both ‘step by step’ and ‘step by step transformation of GT’ datasets, suggests fewer details in its content. Despite this, the performance of the data directly generated by GPT-4 frequently ranks among the best of all GPT-4 answer variants. Meanwhile, other data generation methods exhibited significantly lower performance than GPT-4’s direct answers on at least two tasks, as highlighted in red in Table 2. This suggests that, given a problem, the data directly generated by an advanced LLM might yield the good training results. We observed the same pattern in Table 6 and Table 7. Artificially defined prompts that add information to the ground truth may lead to the generation of target responses that do not conform to the LLM’s linguistic or logical preferences.
Experiment 2: The impact of familiarity
@@
GPT-4 Answer with Lower Perplexity Mistral 0.600 2.20 0.303 2.10 0.722 6.15
@@
GPT-4 Answer with Lower Perplexity Llama2 0.424 2.21 0.153 2.36 0.655 5.11
@@
Claude Answer with Lower Perplexity Mistral 0.586 2.36 0.277 2.24 0.714 4.02
Claude Answer with Higher Perplexity 0.494 5.96 0.218 8.98 0.7 10.63
Claude Answer with Lower Perplexity Llama2 0.433 2.04 0.137 2.33 0.678 3.51
Claude Answer with Higher Perplexity 0.326 3.36 0.121 5.51 0.619 5.69
Table 3: GPT-4/Claude 3.5: Answers with Lower Perplexity vs. Higher Perplexity. $n_{\text{train}}=1000$.
This section introduces a method for assessing the influence of style ‘familiarity’. We direct advanced LLMs to produce several varient of responses to the same query, grouping them into two distinct datasets based on their perplexity levels—high and low. We then evaluate and compare the learning outcomes using these two sets. One concern of this experiment is that advanced LLMs may produce answers with different level of details. To eliminate this factor, we first let advanced LLMs to generate one answer and then direct it to paraphrase the generated answer. In this way, we can create responses that vary in perplexity but maintain the same semantic meaning.
After training on those two sets of training data, we examine their performance, shown in Table 3. As seen, training on the higher-perplexity training set consistently shows slightly lower performance than the lower perplexity counterpart, particularly noticeable in the in-domain performance on two math datasets. For example, when training the Mistral model on the Math Algebra dataset, the GPT-4 answer with lower perplexity set achieves a 30% accuracy rate, while the GPT-4 answer higher perplexity set only reached 23%. Since they are only differed in the style of the text, this result demonstrated that the familiarity of LLMs with the stylistic aspects of target responses significantly affects training outcomes.
@@
Discussion: Why are LLMs often familiar with responses generated from other LLMs? Huh et al. (2024) propose the Platonic Representation Hypothesis, suggesting that different AI models exhibit convergent behavior. This hypothesis supports our findings that LLMs possess intrinsic familiarity with both themselves and others. Additionally, the familiarity may stem from overlapping training data used during pre-training stages, as many LLMs, including GPT-4, share similar datasets, influencing their output styles and content.
@@ 4.2 Achieving high familiarity by exploring the target LLM generated responses
Method Model Type GSM8K Math Algebra ECQA
Groundtruth Mistral 0.434 0.206 0.68
Correct predicted sample + groundtruth 0.491 0.227 0.578
Correct preidcted sample only 0.449 0.257 0.586
Style Transfered Groundtruth 0.51 19.6 0.69
Percentage of training data used in Style Transfered Groundtruth 73.7% 24% 91.2%
Table 4: Performance Comparison when not using advanced LLMs. $n_{\text{train}}=1000$.
In the experiments described above, we validated two hypotheses. All the studies mentioned use advanced LLMs to generate responses, chosen for its advanced capabilities to directly produce correct answers and reliably follow diverse instructions. However, there is a concern regarding the "familiarity" factor: would this argument still hold without the involvement of an advanced LLM? If valid, the argument should apply to some extent even without using advanced LLMs.
This section seeks to develop a method that enhances the “familiarity” of responses without requiring more advanced LLMs than the target LLM.
Use the same model for train and groundtruth style transfer: We devised a method to rewrite the ground truth data in a style that resembles language model-generated responses using the Mistral-7B-Instruct-V2 model. The experiment results are shown in Table 4. Initially, we posed questions from the training dataset and recorded the answers that were correctly predicted by Mistral. We documented two such correct predictions. These two correct responses were then used as in-context examples to guide Mistral on how to rewrite ground truth data. Specifically, we used two recorded predictions as example target responses, their corresponding questions as example questions, and the related ground truth as the example ground truth. The in-context prompt informed Mistral that its task was to rewrite the ground truth in its own styles. For more details, please refer to Figure 6.
@@
We noticed that even when provided with the ground truth, the Mistral model struggles to successfully rewrite answers for mathematical tasks. Therefore, we ran the rewritten math answers through our evaluation script and found that only 73.7% of the GSM8K answers, 91.2% of the ECQA answers and 24% of the Math Algebra answers passed the script. We only used the answers that passed the script for training.
@@
This experiment was conducted solely with Mistral and not with the Llama model, due to our observations that Llama2-13b-chat lacks sufficient in-context following capabilities required for successful ground truth style transformation. We show a failure example in AppendixD. As a comparative experiment, we created the ‘correct predicted samples only’ dataset. For each sample in this dataset, we had the Mistral model perform zero-shot generation 5 times, and using an evaluation filter, we randomly selected one correct prediction that passed the filter to be paired with the original input as a training sample. If the model failed to generate a correct answer in all 5 attempts, we did not include that data in the training dataset. The construction method for the ‘correct predicted samples’ and ‘correct predicted samples only’ datasets is similar. The only difference is that if the sample fails to produce a correct answer, we include the ground truth in the dataset as the correct answer.
@@
As shown in Table 3, the Mistral model trained with style-transferred ground truth data outperformed the model trained with original ground truth data and other methods in all cases except for the Math Algebra dataset, where only 24% of the training data was used. To fairly compare the training effectiveness of the ‘style transferred ground truth’ and the original ground truth, we recorded the IDs of the data used by the ‘style transferred ground truth’. We identified a corresponding set of 240 original ground truth data points for comparison. We found that training with only these 240 data points resulted in a 17.4% accuracy using the original ground truth, which is lower than the 19.6% achieved by the ‘style transferred ground truth’. This demonstrates that data processed using the ‘ground truth style transfer’ method yields better training outcomes than data labeled by humans.
@@
Interestingly, on the ECQA dataset, models trained with correct answers generated by Mistral in a zero-shot fashion performed worse than those trained directly with ground truth. By reading the ‘correct predicted samples only’ dataset, we found that many instances classified as correct by the evaluation filter did not result from sound reasoning by the model. Therefore, the effectiveness of the self-training method depends significantly on whether the evaluation filter used to select correct predictions can truly identify accurate responses.
Method Model Type GSM8K Math Algebra ECQA MBPP HumanEval Avg Perplexity Avg Token Length
GPT-4 Answer Directly Mistra 0.597 0.302 0.722 0.354 0.365 3.81 164.642
Minimum Change on Mistral Predictions 0.562 0.314 0.699 0.354 0.409 2.47 133.944
Minimum Change on LlaMA Predictions 0.547 0.297 0.709 0.364 0.427 3.51 132.323
Average Token Length for Mistral Initial Predictions 152.993
GPT-4 Answer Directly Llama 0.428 0.150 0.656 0.2 0.158 3.58 167.469
Minimum Change on LlaMA Predictions 0.433 0.166 0.649 0.2 0.213 2.75 132.323
Minimum Change on Mistral Predictions 0.402 0.161 0.647 0.218 0.183 3.32 133.945
Average Token Length for Llama Initial Predictions 165.793
Table 5: Comparing the experimental results of GPT4 and minimum change. $n_{\text{train}}=1000$.
@@ 5 A hybrid approach: creating responses using “minimal change” principal
We call this approach “minimal change” and its construction process is as follows: For a given question, we first request a response from the target LLM, such as Llama2 or Mistral. We then instruct GPT-4 to copy the correct portions of the prediction and revise any erroneous parts of the response. When correcting errors, GPT-4 is required to make minimal changes to accurately adjust the answer. Note that this construction process ensures that the target LLM is “familiar” with the response as the response is largely generated from itself. The “minimal change” principal, on the other hand, minimizes the influence from the GPT-4.
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ‘Minimum Change Prompt and Method’ section of the Appendix C.
Results from our ‘Minimum Change’ experiments are presented in Table 5. The data consistently show lower perplexity compared to ground truth data across all datasets, indicating a high level of model familiarity with LLM-generated data. Furthermore, performance across various setups was not only similar (with in-domain performance differences within 5% on most datasets) but also comparable to models trained with direct GPT-4 responses. Although the data was minimally modified based solely on outputs from Mistral/Llama2, theoretically, the logic and quality of the modified answers, as well as the level of detail, should be inferior to answers directly generated by GPT-4. To verify whether the minimum change approach minimally added details, we measured the overall average token length. We first calculated the average token length for each dataset—GSM8K, Math Algebra, ECQA, MBPP, and HumanEval—and then averaged these results to obtain a composite mean. We found that the target responses from the minimum change were not only significantly shorter than those from ‘direct GPT-4 answers’ but also shorter than the initial predictions from Mistral/Llama2. Surprisingly, on the HumanEval and Math Algebra tasks, the performance of models trained with minimally changed data even surpassed those trained with data generated by GPT-4. On other datasets, the performance of the minimum change data was comparable to that of GPT-4. This underscores the effectiveness of training LLMs on data distributions familiar to them, demonstrating that minimal changes to maintain the original structure can be as effective as more detailed alterations. For more experiment details, see Appendix B.
@@ 6 Related Works
Training on synthetic data has been widely employed to enhance model performance, particularly when training data is scarce. Various studies have focused on generating additional data to augment existing datasets Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Piedboeuf and Langlais (2023a); Agrawal et al. (2023); Piedboeuf and Langlais (2023b); Yu et al. (2024) or to create large-scale instruction-tuning datasets when high-quality human annotations are limited Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). In contrast, our work argues that for tasks requiring chain-of-thought reasoning, given the same question-answer pairs from the training dataset, the equivalent answers generated by LLMs are theoretically more effective than human-written ground truth due to familiarity.
Previous works discuss various factors that improve training performance in reasoning, including adding complexity Xu et al. (2023), increasing diversity Luo et al. (2023), incorporating step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023), adding details Zhang et al. (2024); Kang et al. (2023); Li et al. (2022), and ensuring correctness Trinh et al. (2024). We argue that familiarity is another important factor that can improve training outcomes, which has been overlooked by previous works.
Several previous works share similarities with aspects of our research. For example, self-training methods like STAR Zelikman et al. (2022), REST Gulcehre et al. (2023), and RESTem Singh et al. (2023) generate datasets by producing samples from LLMs, filtering out mispredicted samples, and subsequently enhancing the models through further training on these filtered samples. The self-distillation method Yang et al. (2024) uses the target language model to rewrite the ground truth labels in its own words. While their self-distillation or rewriting processes are similar to the methods we employ in Section 4.2, they do not explore using models other than the target model. Our hypothesis suggests that data generated by any LLM is inherently familiar to the target LLM, indicating that more powerful models, such as GPT-4, could be used to generate rewritten labels instead of relying solely on the smaller target model being fine-tuned. Furthermore, these works do not identify the impact of the familiarity.
Similar to our minimum-change approach, the works on learning from mistakes Singh et al. (2023) and process supervision Lightman et al. (2023); Luo et al. (2024) also involve steps of correcting a target LLM generated data. However, their overall process is different from our minimum-change. Moreover, they do not recognize the potential contribution of the familiarity in our approach and our work is complimentary to their discoveries. While they focus on learning from mistakes, our approach centers on maximizing familiarity.
Perplexity has been used in previous studies primarily for data selection and filtering. Gonen et al. (2022) use perplexity to select prompts, demonstrating that lower perplexity prompts lead to better zero-shot QA performance on frozen models. Others have used perplexity to select pre-training data De la Rosa et al. (2022), detect AI-generated code Xu and Sheng (2024), and predict the synthetic generalization of language models Hu et al. (2020). Perplexity has also been employed to detect language model attacks Alon and Kamfonas (2023) and to select instruction tuning data Li et al. (2024); Mekala et al. (2024), where it reflects data difficulty. In contrast, we use perplexity to measure familiarity. Our focus is on how different outputs $y$, given the same input $x$, can affect training effectiveness due to variations in perplexity. We use perplexity to guide the construction of target responses that are more familiar to the model.